BUSINESS TRANSACTIONS for the Month of December.

1) 12-1 The Shark Corp Received $220,000 from SJ Sharkie in exchange for 5,000 shares of Common Stock.

2) 12-2 The Shark Corp Borrowed $45,000 from Wells Fargo Bank and signed a Note.

3) 12-3 The Shark Corp purchased Office Furniture for $120,000(5 year life), on account from IKEA..

4) 12-4 The Shark Corp purchased Computer Equipment for $60,000(5 year life), on account from Office Depot

5) 12-5 The Shark Corp paid $18,000 with Check#101 for Corporate Insurance for the next 12 months.(Dec-Nov)

6) 12-10 The Shark Corp paid $20,000 with Check#102 for December Rent.

7) 12-11 The Shark Corp provided a mascot Seminar for the NIKE Corp and Billed Nike $120,000 on account.

8) 12-15 The Shark Corp received the December Utility Bill from PGE in the amount of $5,000 to be paid in Jan.

9) 12-16 The Shark Corp received the December Telephone Bill from SBC for $6,000 to be paid Later.

10) 12-17 The Shark Corp received a $70,000 Check from Nike Corp on account for services previously provided.

11) 12-20 The Shark Corp paid $10,000 with check #103 to Mark's Marketing for Professional Fees Expense.

12) 12-25 The Shark Corp paid $44,000 with Check #104 for December Salary expenses.

13) 12-27 The Shark Corp received a $25,000 Check from Buick Corp for services to be provided in January.

14) 12-28 The Shark Corp Paid $120,000 with Check#105 to IKEA Corp for the Office Furniture purchased above.

15) 12-29 The Shark Corp Paid $6,000 with Check #106 to SBC on account for the Telephone Bill.

16) 12-30 The Shark Corp Paid $12,000 with Check#107 to Mr. SJ Sharkie for Dividends to the owner.

17) 12-30 The Shark Corp paid $3,000 with check #108 to WF Bank on the note, $1,000 was for Interest Expense.

18) 12-31 The Shark Corp recorded one Month of Insurance expense from the Prepaid Insurance account.

19) 12-31 The Shark Corp recorded one Month of Depreciation Expense on the Furniture and the Computer Equip.

20) 12-31 The Shark Corp recorded Income Tax Payable and Income Tax Expense for December, which is 40% of Net Income (Taxes to be discussed in class)

____________________________________________________________________________________________________________________________________________________

Above are the transactions I need to journalize. Below, I have answers for all of them except #19 and #20, as I am confused as to how you get the value of one-month depreciation expenses for furniture and comp. equipment. Please explain how to do #19, as well as 20, and correct my answers if they're wrong.

Date                   Account Description               Debit               Credit

1-Dec                   Cash                                       220,000  

                               Common Stock                                           5,000

2-Dec                   Cash                                       45,000  

                               Notes Payable - Wells Fargo                       45,000

3-Dec                   Office Furniture                       120,000  

                               Accounts Payable - IKEA                           120,000

4-Dec                   Computer Equipment               60,000  

                               Accounts Payable - Office Depot               60,000

5-Dec                   Prepaid Insurance                   18,000  

                               Cash                                                           18,000

10-Dec               Rent Expense                           20,000  

                               Cash                                                           20,000

11-Dec                   Accounts Receivable               120,000  

                               Service Revenue                                           120,000

15-Dec               Utility Expense                       5,000  

                               Utilities Payable                                           5,000

16-Dec               Telephone Bill Expenses           6,000  

                               Telephone Bill Payable                               6,000

17-Dec               Cash                                       70,000  

                               Accounts Receivable                                   70,000

20-Dec               Professional Fees Expense   10,000  

                               Cash                                                           10,000

25-Dec               Salary Expenses                       44,000  

                               Cash                                                            44,000

27-Dec               Cash                                       25,000  

                               Unearned Service Revenue                       25,000

28-Dec               Accounts Payable - IKEA       120,000  

                               Cash                                                           120,000

29-Dec               Telephone Bill Payable           6,000  

                               Cash                                                               6,000

30-Dec               Dividends                               12,000  

                               Cash                                                               12,000

30-Dec               Notes Payable - Wells Fargo   2,000  

                               Interest Expense                                               1,000  

                               Cash                                                               3,000

31-Dec       Insurance Expense ($18,000/12m)   1,500  

                               Prepaid Insurance                                           1,500

31-Dec   Depreciation Expenses                           ?

                   Accumulated Depreciation - Furniture                           ?

                   Accumulated Depreciation - Office Equipment               ?
Answer #1

Let us first calculate depreciation expense

1. Furniture

Total value $ 120,000
Life 5 years
Depreciation per year = 120,000 / 5 = $ 24,000
Depreciation per month = 24,000 / 12 = $ 2,000

2. Computer equipment

Total value $ 60,000
Life 5 years
Depreciation per year = 60,000 / 5 = $ 12,000
Depreciation per month = 12,000 / 12 = $ 1,000

Journal entry

Depreciation Expense 3,000
Accumulated Depreciation - Furniture 2,000
Accumulated Depreciation - Office Equipment 1,000

For 20,

Tax Expense 11,800
Tax Payable 11,800

Note - Credit amount is incorrect in the first entry.

Revenue $         120,000
Insurance Expense $              1,500
Rent Expense $           20,000
Utility Expense $              5,000
Telephone Expense $              6,000
Professional Fee $           10,000
Salary $           44,000
Interest Expense $              1,000
Depreciation $              3,000
Total Expenses $           90,500
Income before taxes $           29,500
Income Tax $           11,800
Income after taxes $           17,700